Problem 6 Intro A bond has an annual coupon rate of 4.3%, a face value of $1,000, a price of $1,196.59, and matures in 10 years. Part 1 Attempt 1/10 for 10 pts. What is the bonds YTM? 4+ decimals Submit
Problem 7 Intro Forever 21 is expected to pay an annual dividend of $3.35 per share in one year, which is then expected to grow by 10% per year. The required rate of return is 14%. Part 1 B Attempt 1/10 for 8 pts. What is the current stock price? No decimals Submit
Answer #1

Sol 1: For Calculating Bond YTM we would use the following formula i.e. YTM= Coupon Payment + (Face Value - Price of Bond) divided by Maturity Period of Bond whole divided by (Face Value + Price of Bond) / 2.

Here Coupon Payment= 4.3% of $1000 = $43; Face Value= $1000; Price of Bond= $1196.59; Maturity Period = 10 years. Inputting these values in the above formula: YTM= 43 + [(1000-1196.59) / 10] / [ (1000+1196.59)/2] = 23.341 / 1098.295 = 2.1252%

Sol 2: To calculate Current stock Price we will use the Dividend or Gordon Growth Model Formula which is = Stock's Next Annual Dividend whole divided by Required Rate of Return Less Dividend Growth Rate. We have Stock Next Annual Dividend: $3.35; Required Rate of Return = 14% or 0.14 and Dividend Growth Rate = 10% or 0.10.

Current Price of Stock = 3.35 / (0.14-0.10) = $ 83.75